# Target Tracking via LiDAR-RADAR Sensor Fusion for Autonomous Racing

### v2

@@ Abstract
High Speed multi-vehicle Autonomous Racing will increase the safety and performance of road-going Autonomous Vehicles. Precise vehicle detection and dynamics estimation from a moving platform is a key requirement for planning and executing complex autonomous overtaking maneuvers. To address this requirement, we have developed a Latency-Aware EKF-based Multi Target Tracking algorithm fusing LiDAR and RADAR measurements. The algorithm explots the different sensor characteristics by explicitly integrating the Range Rate in the EKF Measurement Function, as well as a-priori knowledge of the racetrack during state prediction. It can handle Out-Of-Sequence Measurements via Reprocessing using a double State and Measurement Buffer, ensuring sensor delay compensation with no information loss. This algorithm has been implemented on Team PoliMOVE’s autonomous racecar, and was proved experimentally by completing a number of fully autonomous overtaking maneuvers at speeds up to 275 km/h.
@@ I INTRODUCTION
Multi-Vehicle Autonomous Racing allows researchers to safely develop and test the software and hardware for fully Autonomous Vehicles (AVs) operating under a variety of limit conditions. The development of algorithms capable of reacting to high-speed unpredictable maneuvers of other vehicles while staying close to their performance bounds will greatly enhance the safety of future commercial AVs. The Indy Autonomous Challenge (IAC), a multi-vehicle racing competition of self-driving cars operated by university researchers, is the optimal proving ground for the development of these technologies.
Vehicle Detection and Tracking (VDT) is a key enabling functionality of AVs, especially for the successful planning and execution of overtaking and defensive racing maneuvers. The key requirements for a VDT algorithm for this application lie in its accuracy and consistency in opponent pose and velocity estimation, especially when close to the EGO vehicle. The algorithm needs to run online while being aware and compensating for the sensor latencies in order to provide the most up-to-date possible estimation of the opponent state.
Figure 1: Visualization of the proposed target tracking algorithm functioning and ground truth evaluation on experimenal data acquired during the 2023 Indy Autonomous Challenge event at the Las Vegas Motor Speedway
In this work, we present a latency-aware Multi-Target Tracking (MTT) algorithm, employing an Extended Kalman Filter (EKF) for state estimation with a reduced-state Constant Velocity and Turn Rate (CVTR) model and a double state and measurement buffer for Out-Of-Sequence Measurement (OOSM) management via Reprocessing. The MTT algorithm performs performs fusion of pre-processed LiDAR and RADAR measurements coming from sensors mounted on the EGO vehicle via explicit integration of the Range Rate measurement into the EKF measurement equation together with a yaw rate exogenous input deriving from the curvature of the track circuit.
The algorithm has been implemented online as it has been deployed on Team PoliMOVE’s Dallara AV-21 ”minerVa” and proved its performance by completing a set of fully autonomous overtaking maneuvers at speeds up to 275 km/h.
@@
  • A latency-aware architecture for an online MTT algorithm with OOSM management with no information.
  • The explicit integration of the RADAR Range Rate measurement in the Tracking EKF.
  • The inclusion of a-priori track knowledge in the opponent motion prediction.
We provide experimental validation of the performance achieved by our methodologies compared to RTK GNSS ground truth data on both real data acquired during the 2023 IAC competition at the Las Vegas Motor Speedway, which include overtaking maneuvers at speeds up to 275 km/h, as well as simulated data.
We present these contributions in the remainder of the paper according to the following structure: in Section II, we analyze the current state of the art and research gaps concerning LiDAR-RADAR Target Tracking for Autonomous Racing. In Section III, we describe in detail the structure and the components of the algorithms we developed, and then in Section IV we validate our claims through the analysis of experimental results. Finally, in Section V, we conclude our work discussing the achieved results and we investigate potential future developments.
@@ II RELATED WORK
Target tracking is the process of recursively the dynamic states of one or more moving objects using sensor measurements. As a critical component of autonomous driving systems, it can be broadly classified into Multi-Target Tracking (MTT) and Extended Target Tracking (ETT) approaches, and then further divided depending on the sensors used and the vehicle dynamic estimator and model employed.
MTT algorithms maintain separate tracks, one for each object, assuming one measurement per object per timestep. Classical MTT frameworks, such as Multiple Hypothesis Tracking (MHT) [1] and Joint Probabilistic Data Association (JPDA) [2], handle data association uncertainties. ETT methods address scenarios where a single extended object generates multiple measurements [3].
For state estimation, the Extended Kalman Filter (EKF) is a popular choice for its computational efficiency and handling of nonlinear measurements [4]. More advanced model-based methods, such as the Unscented Kalman Filter (UKF) and Particle Filters (PF), have been developed for stronger nonlinearities or non-Gaussian distributions [5]. Although computationally intensive, these filters can offer significant accuracy improvements over EKF in challenging scenarios.
The chosen motion model significantly impacts tracking perfromance for model-based estimators like EKF and UKF. Simple models like constant velocity (CV) and constant acceleration (CA) serve as basic assumptions for target dynamics [6]. However, vehicles often perform complex maneuvers better represented by nonlinear models, such as Constant Turn Rate And Velocity (CTRV) and Constant Turn Rate And Acceleration (CTRA)[7]. Advanced modeling techniques, like the Interacting Multiple Model (IMM) estimator, combine multiple motion models to adapt dynamically to changing vehicle behavior [8], although at the cost of increased computational and implementation effort.
Sensor fusion [9], particularly combining LiDAR and RADAR, leverages complementary sensor strengths to improve tracking robustness and accuracy [10]. LiDAR offers precise spatial positioning, while RADAR provides larger range and direct Doppler range-rate measurements. [11] and [12] demonstrated superior vehicle tracking performance by integrating LiDAR and RADAR data, although they lack the explicit integration of Range Rate Measurements. Fusion frameworks typically manage sensor latency and unsynchronized measurements via buffering and augmented filtering techniques [10].
@@
III METHODOLOGIES
In this section, we present our approach for Target Tracking via LiDAR-RADAR sensor fusion for autonomous racing. We developed a Multi Target Tracking (MTT) algorithm using an Extended Kalman Filter (EKF) for track dynamics estimation enhanced with a measurement buffering system that compensates for sensor latency and manages Out-of-Sequence Measurements (OOSM). We model the vehicle dynamics using a Constant Velocity Turn Rate (CVTR) model and introduce a Yaw Rate Exogenous Input to further refine state estimation accuracy. A Finite State Machine (FSM) manages track initialization and termination, while Data Association is handled through a Global Nearest Neighbor (GNN) search.
We propose an algorithm that works under the availability of the following signals:
  • A smooth 2D absolute EGO vehicle pose and velocity estimation e.g. from an INS/GNSS receiver
  • A set of LiDAR object 2D position measurements e.g. from a clustering algorithm [13]
  • A set of RADAR object 2D positions and Range Rate Measurements
All of the aforementioned signals come from sensors mounted on the moving sensing platform (EGO Vehicle), although the methods can be expanded to use signals coming from external sources like V2V communication. The sensors are supposed to be calibrated w.r.t. a common reference frame fixed on the EGO vehicle. We assume the sensor measurement delay to be known, even if time-varying. If no reliable delay measure is available, an average constant value should be identified experimentally via suitable calibration procedures.
@@ III METHODOLOGIES
The proposed Target Tracking algorithm structure is shown in Figure 2, and is implemented as a ROS2 node. All the elements composing the algorithm will be explained in more detail within the following subsections.
III-A MTT cycle with Variable Time Step
@@
In order to operate online, we implemented the algorithm as a fixed-rate ROS2 node with a Measurement Buffer $B_{M}$. The main algorithm iterates at a fixed frequency of 33 Hz, higher than the frequency of both sensors. Incoming measurements between Target Tracking iterations are accumulated in $B_{M}$, which is implemented as an Ordered Map with the measurement timestamp $t_{meas}$ as the map key.
@@
At each algorithm iteration, we perform a full Multi Target Tracking Update Cycle, composed of EKF Prediction, Data Association, Track State Management and EKF Correction, for every measurement in $B_{M}$, independently of its kind (LiDAR or RADAR), as from Figure 2. Every measurement is removed from the buffer after its processing, and the next measurement is processed. The iteration ends when the buffer is empty. If the buffer was already empty at the beginning of the iteration, we still perform the Target Tracking cycle, even if only EKF Prediction actually takes place, as the other stages aren’t defined without new measurements.
@@
After processing a measurement, the MTT timestamp $t_{MTT}$ is updated to the timestamp of the most recent processed measurement, $t_{meas}^{old}$. If no measurement has been processed in the current iteration, $t_{MTT}$ is incremented by the time elapsed since the last iteration. The time step $T$ used for the EKF model discretization is computed as $T=t_{meas}-t_{MTT}^{old}$, with $t_{MTT}^{old}$ denoting the value of $t_{MTT}$ before processing the current measurement. To compensate for the sensor delay, the output of the MTT algorithm is always subject to a last EKF prediction by a time step equal to $T_{out}=t_{curr}-t_{MTT}$ with $t_{curr}$ being the current time. The EKF state is not changed during this last prediction.
III-B Out-of-Sequence Measurements Management
@@
If the measurement delay is large or varies greatly between the sensors, it may happen to have $T<0$. This is the definition of an Out-of-Sequence Measurement (OOSM), as we need to process a measurement older than the current state estimate.
We implemented a Reprocessing strategy to handle OOSMs, which involves the use of an additional buffer, the Past States and Measurements Buffer $B_{PSM}$, again implemented as an Ordered Map. After each MTT cycle update, we add a snapshot of the current Track EKF states and the corresponding input measurements (if any) to $B_{PSM}$, with $t_{meas}=t_{MTT}$ as the key.
@@
When an OOSM occurs, we traverse $B_{PSM}$ from the last element until we find a states/measurements pair that is older than the OOSM timestamp. In the process, we add all the encountered measurements to $B_{M}$. We then roll back the MTT algorithm’s state to the snapshot from $B_{PSM}$ that precedes the OOSM, and proceed to process all the measurements added to $B_{M}$, including the OOSM that triggered the Reprocessing, following the standard MTT cycle.
III-C EKF Update, State and Input
@@
$X(k+1)$ $=\mathbf{F}(X_{k},u_{k})+Q$ (1)
$Y(k)$ $=\mathbf{H}(X_{k})+R$
As the EKF state update function $\mathbf{F}$, we employ a reduced-state version of the Constant Velocity and Turn Rate (CVTR) model in its polar form [7]. This model describes the motion of a kinematic point along a circular path, with constant linear and angular velocity, and it is widely employed in Target Tracking literature. It is described in continuous time form as
@@
where the state vector $X(t)=[x,y,v,\theta]\in\mathbb{R}^{4}$ describes the target pose and linear velocity in an inertial, earth-fixed reference frame. The yaw rate $\omega$ is defined as an exogenous input $u\in\mathbb{R}$.
@@
The CVTR model allows for the exact discretization, which holds for any time step, enabling us to vary the time-step at runtime to accommodate non-constant sensor sampling rate and delay, as described by
@@
$\mathbf{X}(k+1)&=\begin{bmatrix}\mbox{$\begin{cases}x_{k}+\frac{2v_{k}}{\omega_{k}}\sin\!\left(\frac{\omega_{k}T}{2}\right)\cos\!\left(\theta_{k}+\frac{\omega_{k}T}{2}\right)&\omega_{k}\neq 0\\[6.0pt] x_{k}+v_{k}\,T\cos\!\left(\theta_{k}\right)&\omega_{k}=0\end{cases}$}\\[8.0pt] \mbox{$\begin{cases}y_{k}-\frac{2v_{k}}{\omega_{k}}\sin\!\left(\frac{\omega_{k}T}{2}\right)\sin\!\left(\theta_{k}+\frac{\omega_{k}T}{2}\right)&\omega_{k}\neq 0\\[6.0pt] y_{k}+v_{k}\,T\sin\!\left(\theta_{k}\right)&\omega_{k}=0\end{cases}$}\\[8.0pt] v_{k}\\[6.0pt] \theta_{k}+\omega_{k}\,T\end{bmatrix}\end{aligned}$
@@
where the Yaw Rate Exogenous Input $\omega$ is computed by multiplying the current track velocity estimate $v(k)=X_{3}$ by the curvature of a reference trajectory parallel to the track centerline. The curvature is computed from the second-order Taylor approximation of the reference trajectory at the point closest to the tracked vehicle pose, as Figure 3 shows, according to:
$u=\omega=v\cdot\rho_{closest}$ (4)
@@
Figure 3: Computation of the Yaw Rate Exogenous Input
@@
The rationale behind the use of this Yaw Rate Exogenous Input comes from three observations: the first one is that a vehicle moving on a racetrack close to its performance limit, even when performing racing maneuvers like overtakes, will still have a very small incidence angle w.r.t. the track centerline, as the majority of its lateral (yaw rate) dynamics will be employed to stay within the track bounds. The second observation lies in the difficulty of estimating a state that is not measured directly but instead is the derivative of another state, again not measured directly and with a non-linear dependence on the measurements. The third observation lies in its minor role in path planning: during overtaking maneuvers the EGO and opponent vehicles will be in close proximity, and therefore the opponent motion prediction will be mostly based on the opponent position, velocity and heading.
III-D EKF Measurement Function

### v3

# Target Tracking via LiDAR-RADAR Fusion for Autonomous Racing

@@ Abstract
High-speed multi-vehicle autonomous racing increases the safety and performance of road-going Autonomous Vehicles. Precise Vehicle Detection and Tracking from a moving platform is a key requirement for planning and executing complex autonomous overtaking maneuvers. To address this requirement, we have developed a latency-aware EKF-based Multi Target Tracking algorithm fusing LiDAR and RADAR measurements. The algorithm exploits the different sensor characteristics by explicitly integrating the range-rate in the EKF measurement function, as well as a priori knowledge of the racetrack during state prediction. It can handle Out-Of-Sequence Measurements via Reprocessing using a double state and measurement buffer, ensuring sensor delay compensation with no information loss. This algorithm has been implemented on Team PoliMOVE’s autonomous racecar, and was validated experimentally by completing a number of fully autonomous overtaking maneuvers at speeds up to 275 km/h.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I INTRODUCTION
Multi-vehicle autonomous racing allows researchers to safely develop and test the software and hardware for Autonomous Vehicles (AVs) operating under a variety of at-the-limit conditions which would be dangerous to replicate with human drivers. The development of algorithms capable of reacting to high-speed unpredictable maneuvers of other vehicles will greatly enhance the safety of future commercial AVs. The Indy Autonomous Challenge (IAC), a multi-vehicle racing competition of self-driving cars operated by university researchers, is the optimal proving ground for the development of these technologies.
Vehicle Detection and Tracking (VDT) is a key enabling functionality of AVs, especially for the successful planning and execution of overtaking and defensive racing maneuvers. The key requirements for a VDT algorithm for this application lie in its accuracy and consistency in opponent pose and velocity estimation. The algorithm needs to run online, be aware of and compensate for sensor latencies in order to provide the most up-to-date possible estimate of the opponent state.
Figure 1: Team PoliMOVE’s Dallara AV-21 (yellow) during an autonomous overtaking maneuver at the Las Vegas Motor Speedway, January 2023. Credits: Indy Autonomous Challenge
In this work, we present a latency-aware Multi-Target Tracking (MTT) algorithm, employing an Extended Kalman Filter (EKF) for state estimation with a reduced-state Constant Velocity and Turn Rate (CVTR) model and a double state and measurement buffer for Out-Of-Sequence Measurement (OOSM) management via Reprocessing. The MTT algorithm performs the fusion of pre-processed LiDAR and RADAR measurements coming from sensors mounted on the ego vehicle via explicit integration of the range-rate measurement into the EKF measurement equation together with a Yaw Rate Exogenous Input derived from the curvature of the track circuit.
The algorithm has been implemented and deployed on Team PoliMOVE’s Dallara AV-21 and has proven its performance by completing a set of fully autonomous overtaking maneuvers at speeds up to 275 km/h.
@@
  • A latency-aware architecture for an online EKF-Based MTT algorithm, allowing OOSM management without information loss.
  • The inclusion of a priori track knowledge in the opponent motion prediction.
  • The explicit integration of the RADAR range-rate measurement in the EKF Measurement Function.
We experimentally validate the performance achieved by our contribution on both real data acquired during the 2023 IAC competition at the Las Vegas Motor Speedway, as well as on simulated data at the Autodromo Nazionale Monza. Ablation studies with respect to the RTK-GNSS ground truth demonstrate the contribution of each algorithmic module to the overall tracking performance.
The remainder of the paper is structured as follows: in Section II, we analyze the current state of the art and research gaps concerning LiDAR-RADAR Target Tracking for Autonomous Racing. In Section III, we describe in detail the structure and the components of the algorithms we developed, and then in Section IV we validate our claims through the analysis of experimental results. Finally, in Section V, we conclude our work discussing the achieved results and we investigate potential future developments.
@@ II RELATED WORK
Target tracking is the process of recursively estimating the dynamic states of one or more moving objects using sensor measurements. As a critical component of autonomous driving systems, it can be broadly classified into Multi-Target Tracking (MTT) and Extended Target Tracking (ETT) approaches, and then further divided depending on the sensors used and the vehicle dynamics estimator and model employed.
MTT algorithms maintain separate tracks, one for each object, assuming one measurement per object per timestep. Classical MTT frameworks, such as Multiple Hypothesis Tracking (MHT) [dayangac2016target] and Joint Probabilistic Data Association (JPDA) [fortmann1983sonar], handle data association uncertainties. ETT methods address scenarios where a single extended object generates multiple measurements [pieroni2024design].
For state estimation, the Extended Kalman Filter (EKF) is a popular choice for its computational efficiency and handling of nonlinear measurements [julier2004unscented]. More advanced model-based methods, such as the Unscented Kalman Filter (UKF) and Particle Filters (PF), have been developed for stronger nonlinearities or non-Gaussian distributions [tian2022comparing]. Although computationally more intensive, these filters can offer significant accuracy improvements over the EKF in challenging scenarios.
The chosen motion model significantly impacts tracking performance for model-based estimators like EKF and UKF. Simple models like constant velocity (CV) and constant acceleration (CA) serve as basic assumptions for target dynamics [li2003survey]. However, vehicles often perform complex maneuvers better represented by nonlinear models, such as Constant turn rate and velocity (CTRV) and constant turn rate and acceleration (CTRA) [roth2014ekf]. Advanced modeling techniques, like the Interacting Multiple Model (IMM) estimator, combine multiple motion models to adapt dynamically to changing vehicle behavior [mazor1998interacting], although at the cost of increased computational and implementation effort.
Sensor fusion [yeong2021sensor], particularly combining LiDAR and RADAR, leverages complementary sensor strengths to improve tracking robustness and accuracy [khaleghi2013multisensor]. LiDAR offers precise spatial positioning, while RADAR provides larger range and direct Doppler range-rate measurements. The authors in [karle2023multi] demonstrate superior vehicle tracking performance by integrating LiDAR and RADAR data, although they do not explicitly integrate range-rate measurements.
@@
III METHODOLOGY
In this section, we present our approach for target tracking via LiDAR-RADAR sensor fusion for autonomous racing. We developed a Multi-Target Tracking (MTT) algorithm using an Extended Kalman Filter (EKF) for track dynamics estimation enhanced with a measurement buffering system that compensates for sensor latency and manages Out-of-Sequence Measurements (OOSM). We model the vehicle dynamics using a Constant Velocity Turn Rate (CVTR) model and introduce a Yaw Rate Exogenous Input to further refine state estimation accuracy. A Finite State Machine (FSM) manages track initialization and termination, while data association is handled through a Global Nearest Neighbor (GNN) search.
We propose an algorithm that assumes availability of the following signals:
  • A smooth 2D absolute ego vehicle pose and velocity estimation e.g., from an INS/GNSS receiver
  • A set of LiDAR-detected objects with 2D position measurements e.g., from a clustering algorithm [cellina2025lidar]
  • A set of RADAR-detected objects with 2D positions and range-rate measurements
All the aforementioned signals come from sensors mounted on the moving sensing platform (ego vehicle), although the methods can be expanded to use signals coming from external sources like V2V communication. The sensors are supposed to be calibrated w.r.t. a common reference frame fixed on the ego vehicle. We assume the sensor measurement delay to be known, even if time-varying. If no reliable delay measurement is available, an average constant value should be identified experimentally via suitable calibration procedures.
@@ III METHODOLOGY
III-A Multi-Target Tracking with OOSM management
@@
In order to operate online, we implemented the algorithm as a fixed-rate ROS2 node with a Measurement Buffer $B_{M}$. The main algorithm iterates at a fixed frequency of 33 Hz, higher than the scan frequency of both sensors. Incoming measurements between Target Tracking iterations are accumulated in $B_{M}$, which is implemented as an Ordered Map with the measurement timestamp $t_{meas}$ as the map key.
@@
At each algorithm iteration, we perform a full MTT Update Cycle, composed of EKF Prediction, Data Association, Track State Management and EKF Correction, for every set of measurements in $B_{M}$, as shown in Figure 2. Every measurement is removed from the buffer after its processing, and the next measurement is processed. The iteration ends when the buffer is empty. If the buffer was already empty at the beginning of the iteration, we still perform the Target Tracking cycle, even if only EKF Prediction actually takes place, as the other stages aren’t defined in the absence of new measurements.
@@
After processing a measurement, the MTT timestamp $t_{MTT}$ is updated to the timestamp of the most recent processed measurement, $t_{meas}^{old}$. If no measurement has been processed in the current iteration, $t_{MTT}$ is incremented by the time elapsed since the last iteration. The time step $T$ used for the EKF model discretization is computed as $T=t_{meas}-t_{MTT}^{old}$, with $t_{MTT}^{old}$ denoting the value of $t_{MTT}$ before processing the current measurement. To compensate for the sensor delay, the output of the MTT algorithm is always subject to a final EKF prediction by a time step equal to $T_{out}=t_{curr}-t_{MTT}$ with $t_{curr}$ being the current time. The EKF state is not changed during this last prediction.
Figure 3: Scheme of the OOSM handling via Reprocessing: while in normal operation (left) the measurements are processed in the order of arrival, when an OOSM occurs (right) the state of the MTT is reset to the last valid measurement before the OOSM timestamp and then the measurements are processed in order.
@@
If the measurement delay is large or varies greatly between the sensors, we may obtain $T<0$. This is the definition of an Out-of-Sequence Measurement (OOSM), as we need to process a measurement older than the current state estimate.
@@
To manage OOSMs, we implemented the Reprocessing strategy shown in Figure 3, which involves the use of an additional buffer, the Past States and Measurements Buffer $B_{PSM}$, again implemented as an Ordered Map. After each MTT cycle update, we add a snapshot of the current Track EKF states and the corresponding input measurements (if any) to $B_{PSM}$, with $t_{meas}=t_{MTT}$ as the key. When an OOSM occurs, we traverse $B_{PSM}$ from the last element until we find a states and measurements pair that is older than the OOSM timestamp. In the process, we add all the encountered measurements to $B_{M}$. We then roll back the MTT state, meaning the state of each track, to the snapshot from $B_{PSM}$ that precedes the OOSM, and proceed to process all the measurements added to $B_{M}$, including the OOSM that triggered the Reprocessing, following the standard MTT cycle.
III-B EKF State, Input and Measurement Function
@@
$X_{k+1}$ $=\mathbf{F}(X_{k},u_{k})+w_{k}$ (1)
$Y_{k}$ $=\mathbf{H}(X_{k})+v_{k}$
with $\mathbf{F}$ being the state transition function, $\mathbf{H}$ the measurement function, and $w_{k}\sim\mathcal{N}(\mathbf{0},\,\mathbf{Q}_{k})$ and $v_{k}\sim\mathcal{N}(\mathbf{0},\,\mathbf{R}_{k})$ representing the process and measurement noise model.
As the EKF state update function $\mathbf{F}$, we employ a reduced-state version of the Constant Velocity and Turn Rate (CVTR) model in its polar form [roth2014ekf]. This model describes the motion of a kinematic point along a circular path, with constant linear and angular velocity, and it is widely employed in Target Tracking literature. It is described in continuous time form as
@@
where the state vector $X(t)=[x,y,v,\theta]\in\mathbb{R}^{4}$ describes the target pose and linear velocity in an inertial, earth-fixed reference frame. We define the turn rate $\omega$, which is the last state of the original CVTR model [roth2014ekf], as an exogenous input for our EKF, $u(t)=\hat{\omega}(t)\in\mathbb{R}$, hence employing a reduced-state CVTR model.
@@
The CVTR model allows for an exact discretization, which holds for any time step, enabling us to vary the time-step at runtime to accommodate non-constant sensor sampling rate and delay, as described by
@@
$\begin{aligned} \mathbf{X}(k+1)&=\begin{bmatrix}\mbox{$\begin{cases}x_{k}+\frac{2v_{k}}{\omega_{k}}\sin\!\left(\frac{\omega_{k}T}{2}\right)\cos\!\left(\theta_{k}+\frac{\omega_{k}T}{2}\right)&\omega_{k}\neq 0\\[6.0pt] x_{k}+v_{k}\,T\cos\!\left(\theta_{k}\right)&\omega_{k}=0\end{cases}$}\\[8.0pt] \mbox{$\begin{cases}y_{k}-\frac{2v_{k}}{\omega_{k}}\sin\!\left(\frac{\omega_{k}T}{2}\right)\sin\!\left(\theta_{k}+\frac{\omega_{k}T}{2}\right)&\omega_{k}\neq 0\\[6.0pt] y_{k}+v_{k}\,T\sin\!\left(\theta_{k}\right)&\omega_{k}=0\end{cases}$}\\[8.0pt] v_{k}\\[6.0pt] \theta_{k}+\omega_{k}\,T\end{bmatrix}\end{aligned}$
@@
where the Yaw Rate Exogenous Input $\omega$ is computed by multiplying the current track velocity estimate $v_{k}$ by the curvature of a reference trajectory parallel to the track centerline. The curvature is computed from the second-order Taylor approximation of the reference trajectory at the point closest to the tracked vehicle pose, as Figure 4 shows, according to:
$u_{k}=\hat{\omega}_{k}=v_{k}\cdot\rho_{closest}(x_{k},y_{k})$ (4)
@@
Figure 4: Computation of the Yaw Rate Exogenous Input using the vehicle estimated velocity $v_{k}$ and the track curvature $\rho_{closest}$
@@
The rationale behind the use of this Yaw Rate Exogenous Input comes from three observations: first, a vehicle moving on a racetrack close to its performance limit, even when performing racing maneuvers like overtakes, will still have a very small incidence angle to the track centerline, as the majority of its lateral dynamics will be employed to stay within the track bounds. The second observation lies in the difficulty of estimating a state that is not measured directly but instead is the derivative of another state, again not measured directly and with a non-linear dependence on the measurements. The third observation lies in its minor role in path planning: during overtaking maneuvers the ego and opponent vehicles will be in close proximity, and therefore the opponent motion prediction will be mostly based on the opponent position, velocity and heading.
III-C EKF Measurement Function
The first step of every MTT cycle is EKF Prediction, in which every track is subject to motion prediction according to (3) and (4). Then, we use the EKF measurement function in order to compute the predicted measurement for every track, enabling data association and EKF correction.